The December 31, 2015, balance sheet of Maria’s Tennis Shop, Inc., showed current assets of $1,025 and current liabilities of $875. The December 31, 2016, balance sheet showed current assets of $1,240 and current liabilities of $915. What was the company's 2016 change in net working capital, or NWC?
Answer #1

Net working capital=Current assets-Current liabilities

Net working capital for 2016=(1240-915)=$325

Net working capital for 2015=(1025-875)=$150

Hence change in Net working capital=(325-150)

=$175.